OpenStudy (anonymous):

Find and simplify the difference quotient f(x+h)-f(x)/h,for the function f(x)=4x^2-x-1

myininaya (myininaya):

\[\frac{(4(x+h)^2-(x+h)-1)-(4x^2-x-1)}{h}\] now multiply combine like terms

OpenStudy (whpalmer4):

Same drill. \[(4*(x+h)^2-(x+h) - 1 - (4x^2-x-1))/h\] \[=(4*(x^2+2xh+h^2)-(x+h)-1-4x^2+x+1)/h\] \[=(4x^2+8xh+4h^2-x-h-1-4x^2+x+1)/h\] \[=(8xh+4h^2-h)/h = (8x+4h-1)/h\] As h->0, that will go to \[8x-1\]

OpenStudy (anonymous):

says that is wrong

myininaya (myininaya):

your question didn't ask for the limit as h->0

OpenStudy (anonymous):

im sorry, it says h cannot equal zero

OpenStudy (whpalmer4):

Sorry, made a mess of the final simplification: \[=(8xh+4h^2-h)/h=8x+4h-1\] (and the limit which wasn't asked for is still 8x-1).